# Deep learning-based sequential data assimilation for chaotic dynamics identifies local instabilities from single state forecasts

### v1

@@ III Numerical results
The $a_{\bm{\theta}}$ operator is trained on the L96 model, and the results will be interpreted and discussed in the context of this model. The L96 model is a one-dimensional model defined over a periodic band of latitude of the Earth atmosphere. Its ordinary differential equations read
@@
where $x_{N_{x}}=x_{0}$, $x_{-1}=x_{N_{x}-1}$, $x_{-2}=x_{N_{x}-2}$, $F=8$, and $N_{\mathrm{x}}=40$ in the basic configuration. The Lyapunov time of this model is $0.60$; it has $13$ positive Lyapunov exponents and a zero Lyapunov exponent as a continuous-in-time autonomous system. Hence, the dimension of ${\mathcal{U}}$ is $N_{\mathrm{u}}=14$.
@@ III.1 Hyperparameters sensitivity analysis
We first carry out a large set of trainings to assess the sensitivity of $a_{\bm{\theta}}$’s performance to its hyperparameters. We chose $N_{\mathrm{iter}}=16$, without any significant gain beyond this value while the numerical cost increases due to a deeper backpropagation. We assume the model to be fully observed with $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$, the identity matrix in $\mathbb{R}^{N_{\mathrm{x}}}$, and the observations to be affected by a white-in-time unbiased Gaussian noise of covariance matrix $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$, for all time steps. This configuration is very often used to benchmark new DA schemes. The ensemble size $N_{\mathrm{e}}$ and the number of filter $N_{\mathrm{f}}$ were selected in a set ranging between $1$ and $40$. The number $N_{\mathrm{b}}$ of residual blocks in the CNN and number of subblocks $N_{\mathrm{sb}}$ in each residual block were both chosen in the set $\llbracket 1,6\rrbracket$. Because the L96 model has short-range correlations in space, we choose a kernel size of $f_{\mathrm{v}}=5$, even though the CNN receptive field is much larger. Overfitting is prevented by an early stopping of the minimization based on the validation score.
@@ III.2 First results and robustness
The test DA runs with the trained $a_{\bm{\theta}}$ are actually all stable in time, yielding an aRMSE significantly below $1$, as expected if DA has any skill over the mere observations.
Unsurprisingly, we found that the larger the hyperparameters $N_{\mathrm{f}},N_{\mathrm{b}},N_{\mathrm{sb}}$, the smaller the test aRMSEs of the resulting DANs, but that $N_{\mathrm{f}}=40$, $N_{\mathrm{b}}=5$, and $N_{\mathrm{sb}}=5$ offer a good compromise for accuracy versus training cost and CNN size. This will be the reference configuration.
One obvious essential drawback of learning $a_{\bm{\theta}}$ is its non-universality. Specifically, $a_{\bm{\theta}}$ depends on the observation setup used in the training dataset. This is a critical research path for end-to-end DA. Although not the aim of the present paper, we nonetheless checked the performance of the trained $a_{\bm{\theta}}$, with $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$ and $\mathbf{R}=\sigma_{\mathrm{y}}^{2}\mathbf{I}_{\mathrm{x}}$ with $\sigma_{\mathrm{y}}=1$, in test DA runs with similar observations but generated with $\sigma_{\mathrm{y}}$ taking value in between $0.1$ and $3.2$. Yet, in all test runs, DAN remains robust with slightly degraded aRMSEs for $\sigma_{\mathrm{y}}<1$ but aRMSEs at least as good for $\sigma_{\mathrm{y}}>1$, compared to a well-tuned EnKF. Well-tuned EnKF always refers here to an EnKF with an ensemble large enough such that localization is unnecessary and relying on the EnKF-N (Bocquet, Raanes, and Hannart, 2015; Raanes, Bocquet, and Carrassi, 2019) to optimally counteract residual sampling errors such that inflation is unnecessary. We also learned a unique $a_{\bm{\theta}}$ from observation networks whose density $N_{\mathrm{y}}/N_{\mathrm{x}}$ is randomly and uniformly chosen in the interval $[0,1]$, and $\sigma_{\mathrm{y}}=1$. This DAN was then tested on several DA runs, each one with a fixed observation density $N_{\mathrm{y}}/N_{\mathrm{x}}$ taking value in the interval $[0.2,1]$. In this configuration, $a_{\bm{\theta}}$ performs almost as well or better than well-tuned EnKFs for $0.650.35<N_{\mathrm{y}}/N_{\mathrm{x}}<$ 0.65 and is suboptimal (compared to the EnKF) but still stable outside of this range. These results already pleasantly suggest that the findings that follow may remain valid well beyond the specifications of observation operators from which $a_{\bm{\theta}}$ was learned.
@@ III.3 One state forecast
Using the reference configuration but with an ensemble size $N_{\mathrm{e}}$ taking values in the set $\llbracket 1,40\rrbracket$, test aRMSEs fluctuate in between $0.19$ and $0.20$. By contrast, a sizable ensemble is, as we recalled in Sec. I, one of the key reason for the success of the EnKF. For quantitative comparison, we checked that 3D-Var scores an aRMSE of $0.40$, the best linear filter, i.e. $a_{\bm{\theta}}$ learned without activation functions scores $0.38$, whereas well-tuned EnKFs with $N_{\mathrm{e}}=20$ and $N_{\mathrm{e}}=40$ score $0.191$ and $0.179$, respectively. Note that the reference $a_{\bm{\theta}}$ but with $N_{\mathrm{f}}=100$ yields an aRMSE of $0.185$, closer to the best EnKF with $N_{\mathrm{e}}=N_{\mathrm{x}}=40$, showing that further improvements are possible even though not the focus of this paper.
However, the key remark is that a single state forecast, $N_{\mathrm{e}}=1$ in $a_{\bm{\theta}}$, is as efficient as using a large ensemble. Furthermore, the need for localization and inflation is completely obviated. We have checked that this is obtained concurrently to a feature collapse in $a_{\bm{\theta}}$,(van Amersfoort et al., 2022) i.e. all channels’ last layer feature maps converge to the same state. It is likely that a better local minimum of the loss could be obtained with complex encoder and decoder (Peyron et al., 2021) and infusing diversity in the NN through Monte Carlo dropouts,(McCabe and Brown, 2021) so as to obtain an $a_{\bm{\theta}}$ leveraging the ensemble. Nonetheless, the local minimum reached in our trainings, yield an accuracy with $N_{\mathrm{e}}=1$ worthy of a well-tuned EnKF. That is why we shall concentrate in the following on interpreting this astonishing result, especially in Sec. IV, using dynamical systems theory.
Consequently, the analysis operator is hereafter learned in the reference configuration but with $N_{\mathrm{e}}=1$.
IV Interpretation and supporting results
In this section, we focus on the main result: a learned DA method with a single state forecast achieves performance on par with a well tuned EnKF. We would like to understand the reason for this performance and what $a_{\bm{\theta}}$ is learning.
IV.1 Kalman gain-based expansion
Towards this goal, we build an incremental classical Kalman update (Kalman, 1960; Ghil and Malanotte-Rizzoli, 1991) that would be a good match to $a_{\bm{\theta}}$ seen as a mathematical map, at least for small analysis increments. The first diagnostic is the mean anomaly generated by $a_{\bm{\theta}}$, i.e. how much $a_{\bm{\theta}}({\mathbf{x}},\mathbf{0})$ deviates from $\mathbf{0}$ on average. It should be small since a vanishing innovation $\bm{\delta}_{k}$ should not yield any state update. Hence, we define the time-dependent normalized scalar anomalies
@@ IV.1 Kalman gain-based expansion
owing to the fact that no state update is needed when the innovation vanishes, and only keeping the leading order term in $\bm{\delta}$. This is an Ansatz of $a_{\bm{\theta}}$ where $\mathbf{K}({\mathbf{x}})\in\mathbb{R}^{N_{\mathrm{x}}\times N_{\mathrm{y}}}$ is meant to stand as a Kalman gain surrogate. With the propagation of a single state, classical sequential DA methods would typically resemble 3D-Var, and the gain would not depend on the forecast state (the first input variable of $a_{\bm{\theta}}$). The 3D-Var equivalent gain would be
$\mathbf{K}_{\mathrm{3D-Var}}=\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\left(\mathbf{R}+\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\right)^{-1},$ (8)
where $\mathbf{R}$ and/or $\mathbf{H}$ might depend on time if the monitoring system does so, and where $\mathbf{B}$ represents some expected climatology of the DA errors.
Interestingly, we also learned $a_{\bm{\theta}}$ replacing Eq. (2a) with the simplified ${\mathbf{E}}_{k}^{\mathrm{a}}={\mathbf{E}}_{k}^{\mathrm{f}}+a_{\bm{\theta}}\left(\mathbf{H}_{k}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}_{k}{\bm{\delta}}_{k}\right)$, whereby losing $a_{\bm{\theta}}$’s ability to extract information from ${\mathbf{E}}_{k}^{\mathrm{f}}$. This yields an aRMSE of $0.38$ in test DA runs, unsurprisingly close to the $0.40$ of our 3D-Var. Hence, learning an optimal constant-in-time $\mathbf{K}$ of an (En)KF(Luk et al., 2024), a configuration subsumed by this specific $a_{\bm{\theta}}$, is significantly suboptimal in this context.
IV.2 Estimation of the Kalman gain and error covariances
Once $a_{\bm{\theta}}$ has been obtained from training, and considering a fixed background state ${\mathbf{x}}$ at a given time step, a large set of innovations $\left\{\bm{\delta}_{j}\right\}_{j=1\ldots N_{\mathrm{p}}}$ are sampled from the observation error statistics: $\bm{\delta}_{j}\sim N(\mathbf{0},\mathbf{R})$. This yields a set of corresponding incremental updates $\left\{\mathbf{a}_{j}=a_{\bm{\theta}}({\mathbf{x}},\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta}_{j})\right\}_{j=1\ldots N_{\mathrm{p}}}$. Since Eq. (7) is only an approximation, $\mathbf{K}({\mathbf{x}})$ is estimated with the least squares problem
@@ IV.2 Estimation of the Kalman gain and error covariances
Next, assuming $\mathbf{R}$ is known, we would like to estimate the analysis $\mathbf{P}^{\mathrm{a}}$ and background $\mathbf{B}$ error covariance matrices, associated to $a_{\bm{\theta}}$ in the Kalman gain interpretation. They depend on ${\mathbf{x}}_{k}$ and hence on $t_{k}$. Given that $\mathbf{K}\in\mathbb{R}^{N_{\mathrm{x}}\times N_{\mathrm{y}}}$, it should not be possible to fully determine either $\mathbf{P}^{\mathrm{a}}$ or $\mathbf{B}$, unless $\mathbf{H}$ is invertible. But we can access error covariances in observation space rather than in state space. Within the best linear unbiased estimator framework, it can be shown that
$\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}=\mathbf{K}\mathbf{R},\qquad\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}=\left(\mathbf{I}_{\mathrm{y}}-\mathbf{H}\mathbf{K}\right)^{-1}\mathbf{H}\mathbf{K}\mathbf{R}.$ (10)
Hence, one can estimate surrogates of $\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}$ and of $\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}$ for the learned $a_{\bm{\theta}}$ from the numerically estimated surrogate $\mathbf{K}$, at any time $t_{k}$. When $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$ and $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$, these boil down to
$\mathbf{P}^{\mathrm{a}}=\mathbf{K},\qquad\mathbf{B}=\left(\mathbf{I}_{\mathrm{x}}-\mathbf{K}\right)^{-1}\mathbf{K}.$ (11)
IV.3 Numerical results
At each $t_{k}$, i.e. over many ${\mathbf{x}}_{k}$ on the attractor of the dynamics, it is possible to estimate, the equivalent $\mathbf{K}({\mathbf{x}}_{k})$, $\mathbf{B}({\mathbf{x}}_{k})$, $\mathbf{P}^{\mathrm{a}}({\mathbf{x}}_{k})$ as seen from $a_{\bm{\theta}}$. For the sake of simplicity, $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$, $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$ are chosen, such that Eq. (11) stands.

### v2

@@ III Numerical results
The $a_{\bm{\theta}}$ operator is trained on the L96 model,(Lorenz and Emanuel, 1998) and the results will be interpreted and discussed in the context of this model. L96 is a one-dimensional model defined over a periodic band of latitude of the Earth atmosphere. Its ordinary differential equations read
@@
with $x_{N_{x}}=x_{0}$, $x_{-1}=x_{N_{x}-1}$, $x_{-2}=x_{N_{x}-2}$, $F=8$, and $N_{\mathrm{x}}=40$ in the basic configuration. The model has a Lyapunov time of $0.60$. It has $13$ positive exponents and, being continuous-in-time and autonomous, it has one zero Lyapunov exponent. Hence, the dimension of its unstable-neutral subspace ${\mathcal{U}}$, is $N_{\mathrm{u}}=14$.
@@ III.1 Hyperparameters sensitivity analysis
We first carry out a large set of trainings to assess the sensitivity of $a_{\bm{\theta}}$’s performance to its hyperparameters. We choose $N_{\mathrm{iter}}=16$, without any significant gain beyond this value while the numerical cost increases due to a deeper backpropagation. We first assume the model to be fully observed with $\mathcal{H}_{k}=\mathbf{I}_{\mathrm{x}}$, the identity matrix in $\mathbb{R}^{N_{\mathrm{x}}}$, and the observations to be affected by a white-in-time unbiased Gaussian noise of covariance matrix $\mathbf{R}_{k}=\mathbf{I}_{\mathrm{x}}$, for all time steps. This configuration is the most widely used to benchmark new DA schemes with L96. The ensemble size $N_{\mathrm{e}}$ and the number of filter $N_{\mathrm{f}}$ were selected in a set ranging between $1$ and $40$. The number $N_{\mathrm{b}}$ of residual blocks in the CNN and number of subblocks $N_{\mathrm{sb}}$ in each residual block were both chosen in the set $\llbracket 1,6\rrbracket$. Because L96 has short-range correlations in space, we choose a kernel size of $f_{\mathrm{v}}=5$, even though the CNN receptive field is much larger.
@@ III.2 First results and robustness
The training dataset size per epoch scales linearly with $N_{\mathrm{r}}$, which is chosen to be $2^{18}$ and further discussed in the supplementary material. The subsequent test DA runs with the trained $a_{\bm{\theta}}$ are actually all stable in time, yielding an aRMSE significantly below $1$, as expected if DA has any skill over the mere observations. Unsurprisingly, we found that the larger the hyperparameters $N_{\mathrm{f}},N_{\mathrm{b}},N_{\mathrm{sb}}$, the smaller the test aRMSEs of the resulting DANs, but that $N_{\mathrm{f}}=40$, $N_{\mathrm{b}}=5$, and $N_{\mathrm{sb}}=5$ offer a good compromise for accuracy versus training cost and CNN size. This will be the reference configuration, which has about $2\times 10^{5}$ trainable parameters.
One obvious essential drawback of learning $a_{\bm{\theta}}$ is its non-universality. Specifically, $a_{\bm{\theta}}$ depends on the observation setup used in the training dataset. This is a critical research path for end-to-end DA. Although not the aim of the present paper, we nonetheless checked the performance of the trained $a_{\bm{\theta}}$, with $\mathcal{H}_{k}=\mathbf{I}_{\mathrm{x}}$ and $\mathbf{R}_{k}=\sigma_{\mathrm{y}}^{2}\mathbf{I}_{\mathrm{x}}$ with $\sigma_{\mathrm{y}}=1$, in test DA runs with similar observations but generated with $\sigma_{\mathrm{y}}$ taking value in between $0.1$ and $3$. Yet, in all test runs, DAN remains robust with slightly degraded aRMSEs for $\sigma_{\mathrm{y}}<1$ but aRMSEs at least as good for $\sigma_{\mathrm{y}}>1$, compared to a well-tuned EnKF. Well-tuned EnKF always refers here to an EnKF with an ensemble large enough such that localization is unnecessary and relying on the EnKF-N (Bocquet, Raanes, and Hannart, 2015; Raanes, Bocquet, and Carrassi, 2019) to optimally counteract residual sampling errors such that inflation is unnecessary.
Testing non-trivial $\mathcal{H}_{k}$, we also learned a single $a_{\bm{\theta}}$ from observation networks whose density $N_{\mathrm{y}}/N_{\mathrm{x}}$ is randomly and uniformly chosen in the interval $[0,1]$ at each $\tau_{k}$, and $\sigma_{\mathrm{y}}=1$. This DAN was then tested on several DA runs, each one with a constant in time observation density $N_{\mathrm{y}}/N_{\mathrm{x}}$ taking value in the interval $[0.2,1]$. In this configuration, $a_{\bm{\theta}}$ performs almost as well or better than well-tuned EnKFs for $0.650.35<N_{\mathrm{y}}/N_{\mathrm{x}}<$ 0.65 and is suboptimal (compared to the EnKF) but still stable outside of this range. These results already pleasantly suggest that these DL-based DA schemes may remain valid well beyond the specifications of observation operators from which $a_{\bm{\theta}}$ was learned. Plots of these experiments and further discussion are provided in the supplementary material.
@@ III.3 One state forecast
Using the reference configuration but with an ensemble size $N_{\mathrm{e}}$ taking values in the set $\llbracket 1,40\rrbracket$, test aRMSEs fluctuate in between $0.19$ and $0.20$. By contrast, a sizable ensemble is, as we recalled in Sec. I, one of the key reason for the success of the EnKF. For quantitative comparison, we checked that 3D-Var scores an aRMSE of $0.40$, that the best linear filter, i.e. $a_{\bm{\theta}}$ learned without activation functions scores $0.384$, whereas well-tuned EnKFs with $N_{\mathrm{e}}=20$ and $N_{\mathrm{e}}=40$ score $0.191$ and $0.179$, respectively. Note that the reference $a_{\bm{\theta}}$ but with $N_{\mathrm{f}}=100$ yields an aRMSE of $0.185$, closer to the best EnKF with $N_{\mathrm{e}}=N_{\mathrm{x}}=40$, showing that further improvements are possible even though not the focus of this paper. These key aRMSE scores are arranged in a table in the supplementary material.
However, the pivotal remark is that a single state forecast, $N_{\mathrm{e}}=1$ in $a_{\bm{\theta}}$, is as efficient as using a large ensemble. Furthermore, the need for localization and inflation is completely obviated. We have checked that this is obtained concurrently to a feature collapse in $a_{\bm{\theta}}$,(van Amersfoort et al., 2022) i.e. all channels’ last layer feature maps converge to the same state. It is likely that a better local minimum of the loss could be obtained with complex encoder and decoder (Peyron et al., 2021) and infusing diversity in the CNN through Monte Carlo dropouts,(McCabe and Brown, 2021) so as to obtain an $a_{\bm{\theta}}$ leveraging the ensemble. Nonetheless, the local minimum reached in our trainings, yield an accuracy with $N_{\mathrm{e}}=1$ worthy of a well-tuned EnKF. That is why we shall concentrate in the following on interpreting this astonishing result for which we shall use, especially in Sec. IV, dynamical systems theory.
Therefore, the analysis operator is hereafter learned in the reference configuration but with $N_{\mathrm{e}}=1$.
IV Interpretation
In this section, we focus on the remarkable finding that a learned DA method with a single state $N_{\mathrm{e}}=1$ forecast achieves performance on par with a well-tuned EnKF. We wish to understand the reason for this performance by investigating what $a_{\bm{\theta}}$ learns. To that end, an innovative expansion of $a_{\bm{\theta}}$ in terms of more familiar DA operators is carried out.
IV.1 Operator expansion of $a_{\bm{\theta}}$
Towards this goal, we look for a classical Kalman update (Kalman, 1960; Ghil and Malanotte-Rizzoli, 1991) that would be a good match to $a_{\bm{\theta}}$ seen as a mathematical map, at least for small analysis increments. The first diagnostic is the mean anomaly generated by $a_{\bm{\theta}}$, i.e. how much $a_{\bm{\theta}}({\mathbf{x}},\mathbf{0})$ deviates from $\mathbf{0}$ on average. It should be small since a vanishing innovation $\bm{\delta}_{k}$ should not yield any state update. Hence, we define the time-dependent normalized scalar anomalies
@@ IV.1 Operator expansion of $a_{\bm{\theta}}$
owing to the fact that no state update is needed when the innovation vanishes, and only keeping the leading order term in $\bm{\delta}$. This is an Ansatz of $a_{\bm{\theta}}$ where $\mathbf{K}({\mathbf{x}})\in\mathbb{R}^{N_{\mathrm{x}}\times N_{\mathrm{y}}}$ is meant to stand as a Kalman gain surrogate. By contrast, with the propagation of a single state, classical sequential DA methods would typically resemble 3D-Var, and the gain would not depend on the forecast state (the first input variable of $a_{\bm{\theta}}$).
Interestingly, we also learned a simplified ${\hat{a}}_{\bm{\theta}}$ replacing Eq. (2a) with ${\mathbf{E}}_{k}^{\mathrm{a}}={\mathbf{E}}_{k}^{\mathrm{f}}+{\hat{a}}_{\bm{\theta}}\left(\mathbf{H}_{k}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}_{k}{\bm{\delta}}_{k}\right)$, whereby losing $a_{\bm{\theta}}$’s ability to extract information from ${\mathbf{E}}_{k}^{\mathrm{f}}$, similarly to 3D-Var. This yields an aRMSE of $0.382$ in test DA runs, unsurprisingly close to the $0.40$ of our 3D-Var. Hence, learning an optimal constant-in-time $\mathbf{K}$ of an (En)KF,(Luk et al., 2024) a configuration subsumed by this specific ${\hat{a}}_{\bm{\theta}}$, is significantly suboptimal in this context.
IV.2 Identifying the operators in this expansion
Once $a_{\bm{\theta}}$ has been obtained from training, and considering a fixed forecast state ${\mathbf{x}}$ at a given time step, a large set of innovations $\left\{\bm{\delta}_{j}\right\}_{j=1,\ldots,N_{\mathrm{p}}}$ are sampled from the observation error statistics: $\bm{\delta}_{j}\sim N(\mathbf{0},\mathbf{R})$. This yields a set of corresponding incremental updates $\left\{\mathbf{a}_{j}=a_{\bm{\theta}}({\mathbf{x}},\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta}_{j})\right\}_{j=1,\ldots,N_{\mathrm{p}}}$. Since Eq. (8) is only an approximation, $\mathbf{K}({\mathbf{x}})$ is estimated with the least squares problem
@@ IV.2 Identifying the operators in this expansion
Next, assuming $\mathbf{R}$ is known, we would like to estimate the analysis error covariance matrix $\mathbf{P}^{\mathrm{a}}$ associated to $a_{\bm{\theta}}$ in the Kalman gain expansion. It depends on ${\mathbf{x}}_{k}$ and hence on $\tau_{k}$. Within the best linear unbiased estimator framework, $\mathbf{K}$ is related to $\mathbf{P}^{\mathrm{a}}$ through $\mathbf{K}=\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}$, so that from Eq. (8):
$a_{\bm{\theta}}({\mathbf{x}},\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta})\approx\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta},$ (10)
which suggests that an expansion in the second variable ${\bm{\zeta}}\in\mathbb{R}^{N_{\mathrm{x}}}$ of $a_{\bm{\theta}}$ yields
$a_{\bm{\theta}}({\mathbf{x}},{\bm{\zeta}})\approx\mathbf{P}^{\mathrm{a}}({\mathbf{x}})\cdot{\bm{\zeta}}.$ (11)
Hence, $\mathbf{P}^{\mathrm{a}}$ can be estimated using Eq. (11) either from a least squares loss similar to Eq. (9) or from the Jacobian of $a_{\bm{\theta}}$ with respect to ${\bm{\zeta}}$ leveraging auto-differentiable DL libraries.
IV.3 What is learned? - Supporting numerical results
At each $\tau_{k}$, i.e. over many ${\mathbf{x}}_{k}$ on the forecast model’s attractor, it is possible to estimate $\mathbf{K}({\mathbf{x}}_{k})$ and $\mathbf{P}^{\mathrm{a}}({\mathbf{x}}_{k})$ from the expansion of $a_{\bm{\theta}}$. For the sake of simplicity, $\mathcal{H}_{k}=\mathbf{I}_{\mathrm{x}}$, $\mathbf{R}_{k}=\mathbf{I}_{\mathrm{x}}$, in which case $\mathbf{P}^{\mathrm{a}}({\mathbf{x}}_{k})=\mathbf{K}({\mathbf{x}}_{k})$.